# VARADE: a Variational-based AutoRegressive model for Anomaly Detection on the Edge

### v1

@@ Abstract
1 Introduction
In any production context, the downtime of a machine due to the sudden breakdown of mechanical, hydraulic, or electrical components leads to severe losses in terms of time and money. For this reason, efforts are spent for the early detection of any irregular behavior of the production line to avoid sudden stops, enable specific preventive maintenance actions, and reduce the environmental impact. This evolution is enabled by transforming traditional production machinery into Cyber-Physical Systems (CPSs), where sensor devices, communication technologies, and data analytics cooperate to manage production failures in advance [1].
In an industrial CPS scenario, the most crucial resource is the availability of data reflecting the different aspects of production. Such data consist of multiple interdependent variables rapidly evolving over time, thus falling under the typical definition of Multivariate Time Series (MTS) [2]. After collection, the time series, originated by heterogeneous sensors and data sources, are integrated through Industrial Internet of Things (IIoT) technologies and made available for anomaly detection, visualization, and analysis [3].
Although extensive research has been carried out on Multivariate Time Series Anomaly Detection (MTSAD), current solutions typically lack the flexibility and scalability that is required for an effective real-time deployment [4, 5]. In most proposed solutions, the raw data are in fact streamed through the IIoT network to a cloud platform [6], where an expert data-driven system is in charge of the anomaly detection stage. This typically results in high latency owed to communication overhead [7]. Unlike Internet of Things (IoT) networks, IIoT networks are characterized by sensors transmitting a massive amount of data that must be processed in real-time [8, 9]. The order of magnitude of the transmitted data can be GB/s for large production plants, making cloud processing impracticable due to the bandwidth requirements and impairments [9]. All such considerations highlight that anomalies should be detected as soon as possible and as close as possible to the monitored CPS, preferably with real-time or near real-time response, rather than on the cloud. This makes edge computing strategic to maintain the overall system functionally safe [2, 10].
@@ 1 Introduction
2 Background and related works
MTSAD scenarios are best addressed with Deep Learning (DL) methodologies, that recently proved to be more effective in tackling complex anomalies in MTS data [11, 12] than traditional anomaly detection methods (e.g., based on clustering or statistical indexes [13]). Nonetheless, most DL-based solutions present major drawbacks in terms of required data transmission and/or high computational cost [5, 14]. On the other hand, light MTSAD models compatible with edge computing are typically based on tiny and scaled Convolutional Neural Networks (CNNs), that need to be trained with huge sets of annotated anomalies [5, 4, 15, 14]. Collecting and annotating such training sets is however unfeasible in most industrial applications [4, 12].
@@ 2 Background and related works
Forecasting-based methods learn to predict a number of time steps leveraging a current context window. Then, they compare the predicted values with the observed ones to identify anomalies [11]. A large number of studies in this group employ autoregressive Long Short-Term Memory (LSTM) networks, a type of recurrent network able to learn long-term time dependencies in multivariate data [16, 17, 18]. A recent work leverages instead a forest of gradient boosted regression trees to detect anomalies in a Digital Twin-driven industrial context, by examining the residuals from the forecasts of an ensemble of weak predictors [19].
Reconstruction-based methods encode the characteristics of a normal time series into a latent representation and learn to reconstruct new data starting from it. The reconstruction error is then exploited to discriminate the anomalous values from the normal ones. The most popular methods in this group are built on top of autoencoders (AEs), encoder-decoder neural networks where the encoder learns a compressed version of the input data, and the decoder learns to recreate the input starting from the encoded representation. Among the others, [20] employed convolutional AEs for anomaly detection in an IoT-inspired environment, and proved that reducing the size, complexity, and training cost of the AE did not lower its ability to identify anomalies.
Figure 1: Architecture of VARADE. Current ($t_{0}$) and past time steps ($t_{-1}\dots t_{-T}$) are processed by a cascade of convolutional layers and a final linear projection. The output is the estimated probability distribution of the next time step, $P(t_{1})$.
Outlier detectors identify anomalies based on their dissimilarity from regular data points in the feature space. Popular edge-friendly examples in this group are based on k-Nearest Neighbors (kNN), identifying anomalous values based on the distance from their neighbours, and Isolation Forest, that uses the number of binary splits necessary for an ensemble of decision trees to isolate the point from the rest of the data [21].
3 Methods
3.1 Proposed solution
@@
Figure 1 illustrates in principle the proposed architecture. The model takes as input the samples at the current ($t_{0}$) and past time steps ($t_{-1}$,…$t_{-T}$, with $T=16$ for visualization purposes), and passes them through a set of convolutional layers with ReLU activations and a linear projection, to finally predict a single future time step, $t_{1}$. The reason behind this architectural choice lies in the consideration that model inference speed of CNNs is commonly limited by memory bandwidth, especially with SIMD implementations for CPUs and CUDA kernels for GPUs [22]. By using convolutions with kernel size and stride of 2, we obtain that the time-dimension is halved at every new layer, leading to very limited memory usage and bandwidth requirements compared to the number of parameters, and hence to faster inference. On the other hand, the number of feature maps is doubled every two layers (see Figure 1), helping the network to learn more complex and abstract features.
@@
Predicting a probability distribution using a neural network is a complex problem, which can be greatly simplified by constraining the distribution to be Gaussian. This approach, known as variational inference [23], leads to a simpler optimization problem where the objective is to find the mean and variance which minimize the loss function (details will follow). The additional advantage of the Gaussian constraint is that the variance can be interpreted as the uncertainty of the prediction: since we expect the model to be more confident in its prediction when the system is operating normally, and less confident when an anomaly is occurring, the variance can be directly used as an anomaly score.
@@
3.2 Derivation of the loss function
@@
$p(y|\mu,\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)$ (1)
@@
$NLL(y|\mu,\sigma^{2})=-\log\left(\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)\right)$ (2)
@@
$NLL(y|\mu,\sigma^{2})=\frac{1}{2}\log(2\pi\sigma^{2})+\frac{(y-\mu)^{2}}{2\sigma^{2}}$ (3)
@@
$NLL(y|\mu,\sigma^{2})=\frac{1}{2}\log(\sigma^{2})+\frac{(y-\mu)^{2}}{2\sigma^{2}}$ (4)
@@
$L_{\text{recon}}=\frac{1}{2}\left(\log(\sigma_{\text{pred},i}^{2})+\frac{(y_{i}-\mu_{\text{pred},i})^{2}}{\sigma_{\text{pred},i}^{2}}\right)$ (5)
@@
$D_{KL}=-\frac{1}{2}\left(1+\log(\sigma_{\text{pred}}^{2})-\mu_{\text{pred}}^{2}-\sigma_{\text{pred}}^{2}\right)$ (6)
@@
$L=L_{\text{recon}}+\lambda D_{KL}$ (7)
@@ 3.3 Baseline solutions
  • Autoregressive Long Short-Term Memory (AR-LSTM). A recurrent architecture featuring 5 LSTM recurrent layers with 256 feature maps each, followed by 2 fully connected layers. The anomaly score is then calculated as the euclidean norm of the difference between predicted and real value, as in many previous works [24, 16, 17, 18]. To find the best configuration for our specific task, we follow the memory-efficient paradigm introduced by [25], and pick a number of layers equal to 5 based on past experiments at a similar window size [26].
  • Gradient Boosted Regression Forest (GBRF). The technique presented in [19], with minor modifications to boost the anomaly detection capabilities: the number of decision trees is increased from 5 to 30 and the dimensionality reduction step is removed. The anomaly score is computed in the same way as for AR-LSTM.
  • Autoencoder (AE). A convolutional autoencoder featuring 6 ResNet blocks [27]. The anomaly score is the euclidean norm of the difference of reconstructed and real value.
  • kNN. Past works show kNN as the best performing nearest neighbour based algorithm for anomaly detection, with anomaly score computed either as the average or the maximum distance from the neighbors [28]. We employ maximum distance with k=5, as it has the best compromise between accuracy and execution time.
  • Isolation Forest. An ensemble of 100 individual decision trees, that isolate each data point into a leaf. The anomaly score of a data point is based on the average path length [21]. As recommended by [21], we use a contamination value of 0.1, which defines the proportion of outliers in the dataset.
3.4 Implementation details
@@
4 Industrial case study
4.1 Kuka anthropomorphic manipulator
@@
Figure 2: Case study setup. The KUKA manipulator is instrumented with different sensors: 7 accelerometers with six axes (one for each joint) and one single-phase power meter. The sensors are connected directly with an embedded board for detecting different classes of anomalies.
4.2 Data stream characterization
The data stream collected from the robotic manipulator consist of 86 channels in total (reported in Table 1), including signals to monitor the action currently performed by the robot (i.e., action ID), its kinematic behavior (Joint Channels) and its extra-functional parameters (Power Channels[29].
@@ 4.2 Data stream characterization
Table 1: Channels description: for each sensor considered, the related variables are listed. The ¡X¿ in the variable name is a label representing the [0,6] index of the joint in which the corresponding IMU sensor is placed on the robot.
@@ 4.4 Experimental results
Table 2: Comparison between the Anomaly Detection models executed in real-time on the two edge processing units.
@@
Figure 3: Inference Frequency vs Accuracy of the anomaly detection models (identified by the marker color). Marker shape represents the adopted edge device (square for Jetson Xavier NX, and triangle for Jetson AGX Orin), while marker size is proportional to power consumption.
@@
5 Conclusions and future works

### v2

@@ Abstract.
Anomaly detection, Edge computing, Deep learning, Cyber-physical systems, Process monitoring.
journalyear: 2024copyright: acmcopyrightconference: 61st ACM/IEEE Design Automation Conference; June 23–27, 2024; San Francisco, CA, USAbooktitle: 61st ACM/IEEE Design Automation Conference (DAC ’24), June 23–27, 2024, San Francisco, CA, USAdoi: 10.1145/3649329.3655691isbn:
Acknowledgements.
The work has been partially supported by PRIN 2022T7YSHJ SMART-IC - Next Generation EU project. This manuscript reflects only the Authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them.
1. Introduction
In any production context, the downtime of a machine due to the sudden breakdown of mechanical, hydraulic, or electrical components leads to severe losses in terms of time and money. For this reason, efforts are spent for the early detection of any irregular behavior of the production line to avoid sudden stops, enable specific preventive maintenance actions, and reduce the environmental impact. This evolution is enabled by transforming traditional production machinery into Cyber-Physical Systems (CPSs), where sensor devices, communication technologies, and data analytics cooperate to manage production failures in advance (Suvarna et al., 2021).
In an industrial CPS scenario, the most crucial resource is the availability of data reflecting the different aspects of production. Such data consist of multiple interdependent variables rapidly evolving over time, thus falling under the typical definition of Multivariate Time Series (MTS) (Li et al., 2023). After collection, the time series, originated by heterogeneous sensors and data sources, are integrated through Industrial Internet of Things (IIoT) technologies and made available for anomaly detection, visualization, and analysis (Yang and Yue, 2023).
Although extensive research has been carried out on Multivariate Time Series Anomaly Detection (MTSAD), current solutions typically lack the flexibility and scalability that is required for an effective real-time deployment (Nain et al., 2022; Yu et al., 2022). In most proposed solutions, the raw data are in fact streamed through the IIoT network to a cloud platform (Kubiak et al., 2022), where an expert data-driven system is in charge of the anomaly detection stage. This typically results in high latency owed to communication overhead (Bacchiani et al., 2022). Unlike Internet of Things (IoT) networks, IIoT networks are characterized by sensors transmitting a massive amount of data that must be processed in real-time (Hu et al., 2019; Qiu et al., 2020). The order of magnitude of the transmitted data can be GB/s for large production plants, making cloud processing impracticable due to the bandwidth requirements and impairments (Qiu et al., 2020). All such considerations highlight that anomalies should be detected as soon as possible and as close as possible to the monitored CPS, preferably with real-time or near real-time response, rather than on the cloud. This makes edge computing strategic to maintain the overall system functionally safe (Li et al., 2023; Vita et al., 2023).
@@ 1. Introduction
2. Background and related works
MTSAD scenarios are best addressed with Deep Learning (DL) methodologies, that recently proved to be more effective in tackling complex anomalies in MTS data (Schmidl et al., 2022; Li and Jung, 2022) than traditional anomaly detection methods (e.g., based on clustering or statistical indexes (Zhou et al., 2022)). Nonetheless, most DL-based solutions present major drawbacks in terms of required data transmission and/or high computational cost (Yu et al., 2022; Liu et al., 2023). On the other hand, light MTSAD models compatible with edge computing are typically based on tiny and scaled Convolutional Neural Networks (CNNs), that need to be trained with huge sets of annotated anomalies (Yu et al., 2022; Nain et al., 2022; Sun et al., 2023; Liu et al., 2023). Collecting and annotating such training sets is however unfeasible in most industrial applications (Nain et al., 2022; Li and Jung, 2022).
@@ 2. Background and related works
Forecasting-based methods learn to predict a number of time steps leveraging a current context window. Then, they compare the predicted values with the observed ones to identify anomalies (Schmidl et al., 2022). A large number of studies in this group employ autoregressive Long Short-Term Memory (LSTM) networks, a type of recurrent network able to learn long-term time dependencies in multivariate data (Cook et al., 2019; Kumar and Agrawal, 2023; Lu et al., 2023). A recent work leverages instead a forest of gradient boosted regression trees to detect anomalies in a Digital Twin-driven industrial context, by examining the residuals from the forecasts of an ensemble of weak predictors (Huang et al., 2021).
Reconstruction-based methods encode the characteristics of a normal time series into a latent representation and learn to reconstruct new data starting from it. The reconstruction error is then exploited to discriminate the anomalous values from the normal ones. The most popular methods in this group are built on top of autoencoders (AEs), encoder-decoder neural networks where the encoder learns a compressed version of the input data, and the decoder learns to recreate the input starting from the encoded representation. Among the others, (Kim et al., 2018) employed convolutional AEs for anomaly detection in an IoT-inspired environment, and proved that reducing the size, complexity, and training cost of the AE did not lower its ability to identify anomalies.
Figure 1. Architecture of VARADE. Current ($t_{0}$) and past time steps ($t_{-1}\dots t_{-T}$) are processed by a cascade of convolutional layers and a final linear projection. The output is the estimated probability distribution of the next time step, $P(t_{1})$.
Outlier detectors identify anomalies based on their dissimilarity from regular data points in the feature space. Popular edge-friendly examples in this group are based on k-Nearest Neighbors (kNN), identifying anomalous values based on the distance from their neighbours, and Isolation Forest, that uses the number of binary splits necessary for an ensemble of decision trees to isolate the point from the rest of the data (Liu et al., 2012).
3. Methods
3.1. Proposed solution
@@
Figure 1 illustrates in principle the proposed architecture. The model takes as input the samples at the current ($t_{0}$) and past time steps ($t_{-1}$,…$t_{-T}$, with $T=16$ for visualization purposes), and passes them through a set of convolutional layers with ReLU activations and a linear projection, to finally predict a single future time step, $t_{1}$. The reason behind this architectural choice lies in the consideration that model inference speed of CNNs is commonly limited by memory bandwidth, especially with SIMD implementations for CPUs and CUDA kernels for GPUs (Ghimire et al., 2022). By using convolutions with kernel size and stride of 2, we obtain that the time-dimension is halved at every new layer, leading to very limited memory usage and bandwidth requirements compared to the number of parameters, and hence to faster inference. On the other hand, the number of feature maps is doubled every two layers (see Figure 1), helping the network to learn more complex and abstract features.
@@
Predicting a probability distribution using a neural network is a complex problem, which can be greatly simplified by constraining the distribution to be Gaussian. This approach, known as variational inference (Blei et al., 2017), leads to a simpler optimization problem where the objective is to find the mean and variance which minimize the loss function (details will follow). The additional advantage of the Gaussian constraint is that the variance can be interpreted as the uncertainty of the prediction: since we expect the model to be more confident in its prediction when the system is operating normally, and less confident when an anomaly is occurring, the variance can be directly used as an anomaly score.
@@
3.2. Derivation of the loss function
@@
(1) $\small p(y|\mu,\sigma^{2})=\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)$
@@
(2) $\small NLL(y|\mu,\sigma^{2})=-\log\left(\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)\right)$
@@
(3) $\small NLL(y|\mu,\sigma^{2})=\frac{1}{2}\log(2\pi\sigma^{2})+\frac{(y-\mu)^{2}}{2\sigma^{2}}$
@@
(4) $\small NLL(y|\mu,\sigma^{2})=\frac{1}{2}\log(\sigma^{2})+\frac{(y-\mu)^{2}}{2\sigma^{2}}$
@@
(5) $\small L_{\text{recon}}=\frac{1}{2}\left(\log(\sigma_{\text{pred},i}^{2})+\frac{(y_{i}-\mu_{\text{pred},i})^{2}}{\sigma_{\text{pred},i}^{2}}\right)$
@@
(6) $\small D_{KL}=-\frac{1}{2}\left(1+\log(\sigma_{\text{pred}}^{2})-\mu_{\text{pred}}^{2}-\sigma_{\text{pred}}^{2}\right)$
@@
(7) $\small L=L_{\text{recon}}+\lambda D_{KL}$
@@ 3.3. Baseline solutions
  • Autoregressive Long Short-Term Memory (AR-LSTM). A recurrent architecture featuring 5 LSTM recurrent layers with 256 feature maps each, followed by 2 fully connected layers. The anomaly score is then calculated as the euclidean norm of the difference between predicted and real value, as in many previous works (Roy et al., 2021; Cook et al., 2019; Kumar and Agrawal, 2023; Lu et al., 2023). To find the best configuration for our specific task, we follow the memory-efficient paradigm introduced by (Sak et al., 2014), and pick a number of layers equal to 5 based on past experiments at a similar window size (Azzouni and Pujolle, 2017).
  • Gradient Boosted Regression Forest (GBRF). The technique presented in (Huang et al., 2021), with minor modifications to boost the anomaly detection capabilities: the number of decision trees is increased from 5 to 30 and the dimensionality reduction step is removed. The anomaly score is computed in the same way as for AR-LSTM.
  • Autoencoder (AE). A convolutional autoencoder featuring 6 ResNet blocks (He et al., 2016). The anomaly score is the euclidean norm of the difference of reconstructed and real value.
  • kNN. Past works show kNN as the best performing nearest neighbour based algorithm for anomaly detection, with anomaly score computed either as the average or the maximum distance from the neighbors (Goldstein and Uchida, 2016). We employ maximum distance with k=5, as it has the best compromise between accuracy and execution time.
  • Isolation Forest. An ensemble of 100 individual decision trees, that isolate each data point into a leaf. The anomaly score of a data point is based on the average path length (Liu et al., 2012). As recommended by (Liu et al., 2012), we use a contamination value of 0.1, which defines the proportion of outliers in the dataset.
3.4. Implementation details
@@
4. Industrial case study
4.1. Kuka anthropomorphic manipulator
@@
Figure 2. Case study setup. The KUKA manipulator is instrumented with different sensors: 7 accelerometers with six axes (one for each joint) and one single-phase power meter. The sensors are connected directly with an embedded board for detecting different classes of anomalies.
4.2. Data stream characterization
The data stream collected from the robotic manipulator consist of 86 channels in total (reported in Table 1), including signals to monitor the action currently performed by the robot (i.e., action ID), its kinematic behavior (Joint Channels) and its extra-functional parameters (Power Channels(Mascolini et al., 2023).
@@ 4.2. Data stream characterization
Table 1. Channels description: for each sensor considered, the related variables are listed. The ¡X¿ in the variable name is a label representing the [0,6] index of the joint in which the corresponding IMU sensor is placed on the robot.
@@ 4.4. Experimental results
Table 2. Comparison between the Anomaly Detection models executed in real-time on the two edge processing units.
@@
Figure 3. Inference Frequency vs Accuracy of the anomaly detection models (identified by the marker color). Marker shape represents the adopted edge device (square for Jetson Xavier NX, and triangle for Jetson AGX Orin), while marker size is proportional to power consumption.
@@
5. Conclusions and future works